# Long Chain-of-Thought for Reasoning Large Language Models

### v1

@@ 4.1.1 Natural Language Deep Reasoning
Traditionally, researchers have sought to adapt natural language for intuitive and free-flowing deep reasoning [594, 781, 204, 435, 749, 548]. Early work by Wei et al. [594] demonstrated that the use of natural language Long CoT significantly enhances the reasoning capabilities of RLLMs. Further, the Natural Program framework [327] allows RLLMs to engage in deeper natural language reasoning by ensuring a more structured and rigorous logical analysis. More recently, CodeI/O [285] has introduced a technique that reorganizes code-based reasoning patterns into natural language formats, further boosting the reasoning potential of RLLMs.
@@ 4.1.2 Structured Language Deep Reasoning
Structured language deep reasoning encompasses various approaches designed to program [70, 330, 483, 418, 132, 599] or symbolic language [425, 104, 321, 264, 654, 424] format for enhanced deep reasoning. In this context, most studies focus on utilizing code to better enhance the mathematical reasoning capabilities [275, 76, 684]. Xu et al. [631] propose a neural-symbol self-training framework guided by the environment, addressing both the scarcity of symbolic data and the limitations of symbolic processing in LLMs. Additionally, Liao et al. [314] present SKIntern, which refines symbolic RLLMs through curriculum learning and linear attenuation, enabling the internalization of symbolic knowledge with fewer examples, reducing computational costs, and accelerating inference. Furthermore, Ranaldi et al. [448] introduce QuaSAR, a CoT variant that directs LLMs to operate at higher abstraction levels through quasi-symbolic reasoning, thus improving natural language reasoning and providing more precise structural representations.
@@ 4.1.3 Latent Space Deep Reasoning
Latent space deep reasoning encompasses techniques designed to enhance the reasoning abilities of LLMs by leveraging operations within continuous latent spaces [481, 100]. These approaches can be categorized into three main paradigms: (1) Reasoning Token-Driven Latent Space Deep Reasoning: Early work [575, 708] introduce the concept of “planning tokens” or “thought tokens” to guide reasoning within latent spaces. Further, Coconut [162] expands on this through the maintenance of multiple alternative reasoning paths, increasing both complexity and efficiency [748, 496]. At the extreme, Heima [469] condenses the entire Long CoT process into a single token, yielding substantial computational savings. (2) Reasoning Vector Driven Latent Space Deep Reasoning: Building on the previous paradigm, LTM [250] conceptualizes the layers of LLMs as “thought blocks” and introduces the concept of “thought vectors” for each layer. This approach allows for the scaling of test-time computations by implicitly performing reasoning within the latent space through recurrent depth. (3) Reasoning Manager Driven Latent Space Deep Reasoning: Inspired by these, Geiping et al. [136] and Saunshi et al. [459] propose a mechanism similar to a continuous reasoning manager, which iteratively governs a trained “recurrent block” as a recurrent “thought block”. This method integrates deeper model layers during reasoning, enhancing performance without needing specialized training data, and even outperforming larger RLLMs. Additionally, ITT [77] leverages the original transformer layer as a recurrent “thought block,” selecting key tokens via adaptive token routing and controlling reasoning depth with residual thinking connections, enabling more efficient processing of critical tokens.
@@ 4.2 Deep Reasoning Learning
Insufficient deep reasoning in RLLMs can significantly degrade performance [542, 587]. As a result, research has focused on improving reasoning through training. Supervised fine-tuning (SFT) [741] stabilizes model outputs by serving as a memory process, while reinforcement learning (RL) enables generalization and self-learning [155, 91]. Recent studies for deep reasoning learning have explored using SFT to imitate advanced reasoning in RLLMs and applying RL to enhance self-improvement in reasoning. As illustrated in Figure 6, this section outlines two key approaches to improve deep reasoning: (1) Deep Reasoning Imitation (§ 4.2.1), which involves learning reasoning from human-annotated or distilled data through SFT, and (2) Deep Reasoning Self-Learning (§ 4.2.2), where models improve reasoning through preference-based RL with implicit rewards. The performance of these methods is shown in Table 2.
@@ 4.2.1 Deep Reasoning Imitation
Deep reasoning in RLLMs can be effectively achieved by mimicking advanced reasoning systems, such as human reasoning [390, 43, 81, 286], advanced RLLMs [155, 41, 670, 262, 72], and Scaling-augmented RLLMs [293, 702, 420, 792]. This approach enables the model to learn complex reasoning patterns and generalize across tasks [657]. Specifically, (1) Imitation from Human: Earlier, Cobbe et al. [95] first propose the deep reasoning imitation paradigm using human examples. ALT [390] improves RLLM reasoning by generating larger datasets of human-annotated logical templates, which fosters deeper reasoning. To enhance diversity, EIT [43] promotes simpler human-generated plans, while LLMs contribute more nuanced reasoning, facilitating collaboration between human input and AI. (2) Imitation from Advanced RLLMs: A body of work utilizes zero-shot prompting to guide large teacher RLLMs in generating reasoning rationale, which is then used to fine-tune smaller RLLMs, marking the beginning of deep reasoning imitation [174, 246, 679]. Additionally, AceMath [347] applies few-shot prompting to distill Long CoT samples from advanced LLMs, followed by multi-stage quality-guided SFT to enhance performance. Chen et al. [76] separate the data synthesis process into planning and reasoning stages, thereby improving reasoning quality. DART-Math [524] effectively distills complex queries requiring deeper reasoning during synthesis, advancing deep reasoning capabilities. (3) Imitation from Scaling-augmented RLLMs: Earlier, Bansal et al. [26] enhance data quality by scaling the sampling size and length, boosting imitation performance. Yang et al. [650] and Zhao et al. [761] further improve data quality by scaling sampling and selecting samples through a reward model. Additionally, Li et al. [293] identify optimal deep reasoning paths through MCTS, advancing imitation effectiveness.
@@ 4.2.2 Deep Reasoning Self-Learning
While simple imitation can yield strong performance, current models still rely heavily on human annotations or outputs from more advanced models for both imitation and distillation. To address this limitation, recent research has focused on enabling more advanced reasoning through techniques like self-play and self-learning [663, 754, 292]. Specifically, self-learning methods can be classified into two paradigms, differentiated by their sampling strategies:
(1) Self-Learning from Direct Sampling: The earliest method, STaR [707], utilizes In-Context Learning (ICL) to sample deep reasoning results and uses the correctness of the final answer as an implicit reward for self-learning [175, 409, 410, 742, 588, 328]. Further, ReST [153] extends this by introducing a Grow-Improve paradigm, where self-generated reasoning is first annotated with rewards and then enhanced via offline RL algorithms. However, these approaches can be fragile, especially when the reward process lacks robustness. Inspired by the Expectation-Maximization (EM) algorithm, Singh et al. [475] propose a method that generates rewards and iteratively optimizes LLMs to achieve the best performance on a validation set, significantly improving robustness. To further strengthen the reward process, Hosseini et al. [179] introduce a method to adapt incorrect solutions, training a verifier to refine the reward process and improve self-learning quality. (2) Self-Learning from Tree Search: Early deep learning methods, such as EXIT [15], combined MCTS with deep neural networks for reinforcement learning, iteratively self-training the network to guide the tree search and enhance reasoning. Building on this, CPO [746] and TPO [316] align each step of Long CoT reasoning with the corresponding tree search path, using Tree of Thoughts (ToT) [668] preference information to support deeper reasoning [665, 203]. Li [302] propose Policy-Guided Tree Search (PGTS), integrating RL with structured tree exploration for more efficient navigation of reasoning paths. Further developments, such as AlphaMath [53], AlphaLLM-CPL [578], and TongGeometry [724], refine MCTS behavior through stepwise trajectory pair extraction and curriculum preference learning, boosting LLM reasoning abilities [431, 295].
@@ 5 Feasible Reflection for Long CoT
Figure 7: The feedback capabilities framework for feasible reflection consists of Overall Feedback and Process Feedback. Overall Feedback includes the Outcome Reward Model (ORM) in a value format, rule extraction for correctness judgment, and overall critic models based on RLLMs. Process Feedback includes the Process Reward Model (PRM) in a value format and step-level critic models, also based on RLLMs.
@@ 5.1.1 Overall Feedback
The overall feedback focuses on providing a global view of the entire process and results, rather than assessing each step individually. This feedback significantly enhances reasoning skills and reward modeling in reinforcement learning for RLLMs. Specifically, as shown in Figure 7 (a), the overall feedback can be categorized into three main sources: Outcome Reward Model, Rule Extraction, and Critic Models Feedback. The performance across these categories is summarized in Table 3.
@@
Although ORM has achieved significant improvements, its accuracy still falls short of 100%, preventing it from outperforming rule-based answer correction feedback [668, 160]. Previous studies, such as STaR [707], ReST [153], and ReFT Trung et al. [531], have demonstrated that feedback based on final answer rewards is more effective than both PRM and ORM in mathematical scenarios [131]. Furthermore, Guo et al. [155] and Xie et al. [622] introduce a multi-stage RL framework that incorporates rule-based rewards, significantly enhancing both output accuracy and length while mitigating reward hacking through simple yet robust rules [24], such as format validation and result verification. In coding scenarios where direct rule-based feedback is difficult, AceCoder [709], O1-Coder [753], and VerMCTS [39] address this challenge by implementing an automated test-case synthesis pipeline, deriving rewards based on program performance [395, 145, 778]. Additionally, Ma et al. [372] propose an automated approach to training a test case generator, which alleviates the scarcity of test cases and demonstrates that increasing the number of test cases correlates with improved reward quality. Moreover, Ma et al. [371] decompose problem-solving into structured subtasks: file localization, function localization, line localization, and code editing generation, and applies multi-viewed rule-based rewards.
Overall Feedback from Critic Models
@@
Research on feedback from Critic Models centers on detecting errors and biases through natural language feedback, also known as self-reflection or self-critique [231, 23, 452, 384, 571, 701]. This method has led to significant improvements across various tasks, particularly in self-correction [600, 772, 137, 121, 752]. Huang et al. [194] contend that traditional LLMs struggle to generate effective feedback without external signals, requiring the development of RLLMs with enhanced feedback capabilities [458]. As a result, many studies leverage RLLMs’ error-identification strengths, often stemming from their pretraining phase, to improve feedback generation and correction [675].
@@
Earlier, McAleese et al. [380] found that training RLLMs to learn self-critique and deep reasoning can further boost performance. Zhang et al. [744] propose a self-contrast mechanism that compares multiple perspectives, identifies differences, and summarizes insights to resolve inconsistencies. However, these methods often offer task-independent feedback. To address this, Hao et al. [161] introduce AutoRace, which tailors evaluation criteria for specific tasks. The Reversal of Thought (RoT) framework [698] introduces a novel paradigm combining reverse reasoning with self-reflection, helping models identify the limits of their knowledge and enhance reasoning efficiency. Furthermore, ACR [779] implements a scoring system for coding tasks, using LLM-as-a-Judge for quality assessment and LLM-as-a-Critic for critiquing low-quality code, improving consistency across benchmarks. Zheng et al. [771] integrate code execution error data and feedback from RLLMs to improve code generation performance. Liu et al. [342] present AGSER, a method using attention-guided self-reflection to address hallucinations by splitting input queries into attentive and non-attentive components. Finally, Saha et al. [456] introduce EvalPlanner, which separates feedback into planning and reasoning components for more streamlined expression using existing RLLMs.
@@ 5.1.2 Process Feedback
Recent studies highlight the significance of feedback in developing effective PRMs for complex reasoning tasks, particularly in a step-level view [88, 303, 366]. (1) Process Annotated PRM Training: Earlier, Lightman et al. [319] demonstrate that training process feedback with human-annotated data (PRM800K) surpasses outcome supervision in creating reliable reward models. However, this approach requires significant human effort. To address this, Wang et al. [567] introduce Math-Shepherd, a dataset that generates step-by-step supervision using a Tree Search-inspired method [52, 700]. Following this, methods like QwQ [517], Skywork-o1 [400], AceMath [347], and PRIME [97] adopt similar techniques to enhance PRM performance. Additionally, Zhang et al. [729] propose entropy regularization to improve model convergence. Rather than focusing solely on the first error step, Full-Step-DPO [636] assigns rewards for the entire reasoning chain, including error steps. VersaPRM [710] extends PRMs across multiple domains, broadening their applicability. Similarly, Gu et al. [148] and Zhang et al. [751] suggest training models with student preferences aligned to teacher preferences, ensuring effective preference distillation. (2) Outcome Annotated PRM Training: Alternative approaches, such as OVM [685], Implicit PRM [699], AutoPSV [350], and DVO [730], leverage outcome supervision or implicit feedback to train PRMs, reducing the need for extensive human-annotated data [627, 456]. UAS [687] incorporates uncertainty-aware value models [187] into feedback predictions. Additionally, Aurora [499] utilizes ensemble prompting strategies and reference answers for reverse verification, training stronger PRMs that better align with the Long CoT data distribution. Furthermore, PAV [462] suggests that rewards should reflect reasoning progress, as measured by changes in the likelihood of producing a correct future response before and after each step. Yang et al. [653], Lee et al. [267], Yoon et al. [683] extend these paradigms to the token level.
@@
Process Feedback from Critic Models
As PRM training remains heavily dependent on manually annotated data, recent research has explored methods for enabling models to generate their natural language feedback to optimize performance [640]. These approaches fall into two primary categories: (1) Model-Driven Feedback Reasoning: Earlier work such as React [669] and Reflexion [471] enhances RLLMs with natural language feedback at each action and reasoning step [130, 89], improving decision-making in diverse tasks. Similarly, Step-DPO [257] uses RLLM to self-verify step-level positive and negative pairs for training through the DPO paradigm, achieving strong performance. Additionally, Sun et al. [492] propose a dynamic error classification framework that adapts based on model outputs, improving performance in mathematical reasoning tasks by addressing specific error patterns in math word problems. Furthermore, Xie et al. [625] and He et al. [168] iteratively apply MCTS to collect preference data, utilizing its forward-looking capabilities to decompose instance-level rewards into more precise step-level signals, thereby enhancing feedback accuracy. However, step-wise feedback often suffers from reliability issues, which can be mitigated by uncertainty quantification [681, 678], improving the reliability of step-wise verification in reward models for mathematical reasoning tasks. Moreover, Fu et al. [123] define the CoT Average Causal Effect (CACE) to capture causal relationships between steps, resulting in a causalized Long CoT where all steps are both correct and comprehensible. (2) Environment-Driven Feedback Reasoning: Given the increasing complexity of large models, there is growing interest in combining prompt-based LLMs with external environments to generate more interpretable and controllable feedback. For example, ORPS [696] and Drori et al. [108] minimize dependence on human annotations by using execution feedback, enabling models to autonomously refine their solutions. Additionally, Shrestha et al. [472] contribute by translating model outputs into Python code, helping to identify logical errors, gain insights into flawed reasoning processes, and guide improvements in mathematical reasoning. Xu et al. [631] integrate reasoning models with an interactive environment, enabling learning in more dynamic scenarios and creating a more generalizable self-learning framework.

### v3

@@ 4.1.1 Natural Language Deep Reasoning
Traditionally, researchers have sought to adapt natural language for intuitive and free-flowing deep reasoning [668, 885, 236, 490, 850, 614, 160]. Early work by Wei et al. [668] demonstrated that the use of natural language Long CoT significantly enhances the reasoning capabilities of RLLMs. Further, the Natural Program framework [366] allows RLLMs to engage in deeper natural language reasoning by ensuring a more structured and rigorous logical analysis. More recently, CodeI/O [324] has introduced a technique that reorganizes code-based reasoning patterns into natural language formats, further boosting the reasoning potential of RLLMs.
@@ 4.1.2 Structured Language Deep Reasoning
Structured language deep reasoning encompasses various approaches designed to program [84, 371, 543, 470, 154, 673, 664] or symbolic language [480, 125, 360, 299, 739, 479, 32] format for enhanced deep reasoning. In this context, most studies focus on utilizing code to better enhance the mathematical reasoning capabilities [313, 91, 775, 73]. Xu et al. [711] propose a neural-symbol self-training framework guided by the environment, addressing both the scarcity of symbolic data and the limitations of symbolic processing in LLMs. Additionally, Liao et al. [353] present SKIntern, which refines symbolic RLLMs through curriculum learning and linear attenuation, enabling the internalization of symbolic knowledge with fewer examples, reducing computational costs, and accelerating inference. Furthermore, Ranaldi et al. [504] introduce QuaSAR, a CoT variant that directs LLMs to operate at higher abstraction levels through quasi-symbolic reasoning, thus improving natural language reasoning and providing more precise structural representations.
@@ 4.1.3 Latent Space Deep Reasoning
Latent space deep reasoning encompasses techniques designed to enhance the reasoning abilities of LLMs by leveraging operations within continuous latent spaces [541, 120, 510]. These approaches can be categorized into three main paradigms: (1) Reasoning Token-Driven Latent Space Deep Reasoning: Early work [646, 805] introduce the concept of “planning tokens” or “thought tokens” to guide reasoning within latent spaces. Further, Coconut [187] expands on this through the maintenance of multiple alternative reasoning paths, increasing both complexity and efficiency [849, 557]. At the extreme, Heima [528] condenses the entire Long CoT process into a single token, yielding substantial computational savings. (2) Reasoning Vector Driven Latent Space Deep Reasoning: Building on the previous paradigm, LTM [284] conceptualizes the layers of LLMs as “thought blocks” and introduces the concept of “thought vectors” for each layer. This approach allows for the scaling of test-time computations by implicitly performing reasoning within the latent space through recurrent depth. (3) Reasoning Manager Driven Latent Space Deep Reasoning: Inspired by these, Geiping et al. [159] and Saunshi et al. [516] propose a mechanism similar to a continuous reasoning manager, which iteratively governs a trained “recurrent block” as a recurrent “thought block”. This method integrates deeper model layers during reasoning, enhancing performance without needing specialized training data, and even outperforming larger RLLMs. Additionally, ITT [93] leverages the original transformer layer as a recurrent “thought block,” selecting key tokens via adaptive token routing and controlling reasoning depth with residual thinking connections, enabling more efficient processing of critical tokens.
@@ 4.2 Deep Reasoning Learning
Insufficient deep reasoning in RLLMs can significantly degrade performance [608, 659]. As a result, research has focused on improving reasoning through training. Supervised fine-tuning (SFT) [840] stabilizes model outputs by serving as a memory process [699], while reinforcement learning (RL) enables generalization and self-learning [179, 109]. Recent studies for deep reasoning learning have explored using SFT to imitate advanced reasoning in RLLMs and applying RL to enhance self-improvement in reasoning. As illustrated in Figure 6, this section outlines two key approaches to improve deep reasoning: (1) Deep Reasoning Imitation (§ 4.2.1), which involves learning reasoning from human-annotated or distilled data through SFT, and (2) Deep Reasoning Self-Learning (§ 4.2.2), where models improve reasoning through preference-based RL with implicit rewards. The performance of these methods is shown in Table 2.
@@ 4.2.1 Deep Reasoning Imitation
Deep reasoning in RLLMs can be effectively achieved by mimicking advanced reasoning systems, such as human reasoning [440, 53, 98, 325], advanced RLLMs [179, 51, 757, 297, 86], and Scaling-augmented RLLMs [332, 796, 473, 899, 35]. This approach enables the model to learn complex reasoning patterns and generalize across tasks [743]. Specifically, (1) Imitation from Human: Earlier, Cobbe et al. [113] first propose the deep reasoning imitation paradigm using human examples. ALT [440] improves RLLM reasoning by generating larger datasets of human-annotated logical templates, which fosters deeper reasoning. To enhance diversity, EIT [53] promotes simpler human-generated plans, while LLMs contribute more nuanced reasoning, facilitating collaboration between human input and AI. (2) Imitation from Advanced RLLMs: A body of work utilizes zero-shot prompting to guide large teacher RLLMs in generating reasoning rationale, which is then used to fine-tune smaller RLLMs, marking the beginning of deep reasoning imitation [199, 280, 769]. Additionally, AceMath [394] applies few-shot prompting to distill Long CoT samples from advanced LLMs, followed by multi-stage quality-guided SFT to enhance performance. Chen et al. [91] separate the data synthesis process into planning and reasoning stages, thereby improving reasoning quality. DART-Math [589] effectively distills complex queries requiring deeper reasoning during synthesis, advancing deep reasoning capabilities. Further, Ahmad et al. [6] propose OpenCodeReasoning, expanding this paradigm to the code scenarios. (3) Imitation from Scaling-augmented RLLMs: Earlier, Bansal et al. [30] enhance data quality by scaling the sampling size and length, boosting imitation performance. Yang et al. [734] and Zhao et al. [863] further improve data quality by scaling sampling and selecting samples through a reward model. Additionally, Li et al. [332] identify optimal deep reasoning paths through MCTS, advancing imitation effectiveness.
@@ 4.2.2 Deep Reasoning Self-Learning
While simple imitation can yield strong performance, current models still rely heavily on human annotations or outputs from more advanced models for both imitation and distillation [396]. To address this limitation, recent research has focused on enabling more advanced reasoning through techniques like self-play and self-learning [750, 855, 331]. Specifically, self-learning methods can be classified into two paradigms, differentiated by their sampling strategies:
(1) Self-Learning from Direct Sampling: The earliest method, STaR [804], utilizes In-Context Learning (ICL) to sample deep reasoning results [523] and uses the correctness of the final answer as an implicit reward for self-learning [200, 460, 461, 841, 660, 368]. Further, ReST [177] extends this by introducing a Grow-Improve paradigm, where self-generated reasoning is first annotated with rewards and then enhanced via offline RL algorithms. However, these approaches can be fragile, especially when the reward process lacks robustness. Inspired by the Expectation-Maximization (EM) algorithm, Singh et al. [534] propose a method that generates rewards and iteratively optimizes LLMs to achieve the best performance on a validation set, significantly improving robustness. To further strengthen the reward process, a series of work introduce a method to adapt incorrect solutions, training a verifier [123, 204] or utilize entropy [645, 829] to select or refine the reward process and improve self-learning quality. (2) Self-Learning from Tree Search: Early deep learning methods, such as EXIT [16], combined MCTS with deep neural networks for reinforcement learning, iteratively self-training the network to guide the tree search and enhance reasoning. Building on this, CPO [846] and TPO [355] align each step of Long CoT reasoning with the corresponding tree search path, using Tree of Thoughts (ToT) [755] preference information to support deeper reasoning [752, 235]. Li [341] propose Policy-Guided Tree Search (PGTS), integrating RL with structured tree exploration for more efficient navigation of reasoning paths. Further developments, such as AlphaMath [63], AlphaLLM-CPL [650], and TongGeometry [821], refine MCTS behavior through stepwise trajectory pair extraction and curriculum preference learning, boosting LLM reasoning abilities [486, 334, 692].
@@ 5 Feasible Reflection for Long CoT
Figure 7: The feedback capabilities framework for feasible reflection consists of Overall Feedback and Process Feedback. Overall Feedback includes the Outcome Reward Model (ORM) in a value format, rule extraction for correctness judgment, and overall RLLMs based on RLLMs. Process Feedback includes the Process Reward Model (PRM) in a value format and step-level RLLMs, also based on RLLMs.
@@ 5.1.1 Overall Feedback
The overall feedback focuses on providing a global view of the entire process and results, rather than assessing each step individually. This feedback significantly enhances reasoning skills and reward modeling in reinforcement learning for RLLMs. Specifically, as shown in Figure 7 (a), the overall feedback can be categorized into three main sources: Outcome Reward Model, Rule Extraction, and RLLMs Feedback. The performance across these categories is summarized in Table 3.
@@
Although ORM has achieved significant improvements, its accuracy still falls short of 100%, preventing it from outperforming rule-based answer correction feedback [755, 185]. Previous studies, such as STaR [804], ReST [177], and ReFT Trung et al. [596], have demonstrated that feedback based on final answer rewards is more effective than both PRM and ORM in mathematical scenarios [153]. Furthermore, Guo et al. [179] and Xie et al. [701] introduce a multi-stage RL framework that incorporates rule-based rewards, significantly enhancing both output accuracy and length while mitigating reward hacking through simple yet robust rules [27], such as format validation and result verification. In coding scenarios where direct rule-based feedback is difficult, OpenCodeInterpreter [875], AceCoder [806], O1-Coder [854], and VerMCTS [49] address this challenge by implementing an automated test-case synthesis pipeline, deriving rewards based on program performance [445, 169, 882]. Additionally, Ma et al. [422] propose an automated approach to training a test case generator, which alleviates the scarcity of test cases and demonstrates that increasing the number of test cases correlates with improved reward quality. Moreover, Ma et al. [421] decompose problem-solving into structured subtasks: file localization, function localization, line localization, and code editing generation, and applies multi-viewed rule-based rewards.
Overall Feedback from RLLMs
@@
Research on feedback from RLLMs centers on detecting errors and biases through natural language feedback, also known as LLM-as-Judge, self-reflection or self-critique [214, 265, 26, 508, 434, 640, 795]. This method has led to significant improvements across various tasks, particularly in self-correction [674, 876, 161, 142, 853]. Huang et al. [223] contend that traditional LLMs struggle to generate effective feedback without external signals, requiring the development of RLLMs with enhanced feedback capabilities [515]. As a result, many studies leverage RLLMs’ error-identification strengths, often stemming from their pretraining phase, to improve feedback generation and correction [764, 34, 36].
@@
Earlier, McAleese et al. [430] found that training RLLMs to learn self-critique and deep reasoning can further boost performance. Zhang et al. [843] propose a self-contrast mechanism that compares multiple perspectives, identifies differences, and summarizes insights to resolve inconsistencies. However, these methods often offer task-independent feedback. To address this, Hao et al. [186] introduce AutoRace, which tailors evaluation criteria for specific tasks. The Reversal of Thought (RoT) framework [792] introduces a novel paradigm combining reverse reasoning with self-reflection, helping models identify the limits of their knowledge and enhance reasoning efficiency. Furthermore, ACR [883] implements a scoring system for coding tasks, using LLM-as-a-Judge for quality assessment and LLM-as-a-Critic for critiquing low-quality code, improving consistency across benchmarks. Zheng et al. [874] integrate code execution error data and feedback from RLLMs to improve code generation performance. Liu et al. [384] present AGSER, a method using attention-guided self-reflection to address hallucinations by splitting input queries into attentive and non-attentive components. Finally, Saha et al. [513] introduce EvalPlanner, which separates feedback into planning and reasoning components for more streamlined expression using existing RLLMs. More comprehensively, Hu et al. [214] outline the complete pipeline, key insights, and practical lessons for training RLLMs to function as judges.
@@ 5.1.2 Process Feedback
Recent studies highlight the significance of feedback in developing effective PRMs for complex reasoning tasks, particularly in a step-level view [106, 342, 415]. (1) Process Annotated PRM Training: Earlier, Lightman et al. [358] demonstrate that training process feedback with human-annotated data (PRM800K) surpasses outcome supervision in creating reliable reward models. However, this approach requires significant human effort. To address this, Wang et al. [634] introduce Math-Shepherd, a dataset that generates step-by-step supervision using a Tree Search-inspired method [62, 794]. Following this, methods like QwQ [581], Skywork-o1 [450], AceMath [394], and PRIME [115] adopt similar techniques to enhance PRM performance. Additionally, Zhang et al. [826] propose entropy regularization to improve model convergence. Rather than focusing solely on the first error step, Full-Step-DPO [716] assigns rewards for the entire reasoning chain, including error steps. VersaPRM [807] extends PRMs across multiple domains, broadening their applicability. Similarly, Gu et al. [172] and Zhang et al. [852] suggest training models with student preferences aligned to teacher preferences, ensuring effective preference distillation. Further, Wang et al. [644] propose VisualPRM400K and expand this paradigm to multimodal scenarios. (2) Outcome Annotated PRM Training: Alternative approaches, such as OVM [776], Implicit PRM [793], AutoPSV [398], and DVO [827], leverage outcome supervision or implicit feedback to train PRMs, reducing the need for extensive human-annotated data [706, 513]. UAS [778] incorporates uncertainty-aware value models [215] into feedback predictions. Additionally, Aurora [561] utilizes ensemble prompting strategies and reference answers for reverse verification, training stronger PRMs that better align with the Long CoT data distribution. Furthermore, PAV [519] suggests that rewards should reflect reasoning progress, as measured by changes in the likelihood of producing a correct future response before and after each step. Yang et al. [738], Lee et al. [302], Yoon et al. [773] extend these paradigms to the token level. Further Liu et al. [389] propose AdaptiveStep for dividing the reasoning trajectory by model confidence. The resulting PRM allows for more accurate token-level evaluations compared to previous PRMs. Moreover, Chen et al. [94] expand these into interactive agent scenarios, allowing for automatically learning reward models from the environment without manual annotation.
@@
Process Feedback from RLLMs
As PRM training remains heavily dependent on manually annotated data, recent research has explored methods for enabling models to generate their natural language feedback to optimize performance [722]. These approaches fall into two primary categories: (1) Model-Driven Feedback Reasoning: Earlier work such as React [756] and Reflexion [530] enhances RLLMs with natural language feedback at each action and reasoning step [152, 107, 77], improving decision-making in diverse tasks. Similarly, Step-DPO [292] uses RLLM to self-verify step-level positive and negative pairs for training through the DPO paradigm, achieving strong performance. Additionally, Sun et al. [553] propose a dynamic error classification framework that adapts based on model outputs, improving performance in mathematical reasoning tasks by addressing specific error patterns in math word problems. Furthermore, Xie et al. [704] and He et al. [193] iteratively apply MCTS to collect preference data, utilizing its forward-looking capabilities to decompose instance-level rewards into more precise step-level signals, thereby enhancing feedback accuracy. However, step-wise feedback often suffers from reliability issues, which can be mitigated by uncertainty quantification [771, 768], improving the reliability of step-wise verification in reward models for mathematical reasoning tasks. Moreover, Fu et al. [145] define the CoT Average Causal Effect (CACE) to capture causal relationships between steps, resulting in a causalized Long CoT where all steps are both correct and comprehensible. (2) Environment-Driven Feedback Reasoning: Given the increasing complexity of large models, there is growing interest in combining prompt-based LLMs with external environments to generate more interpretable and controllable feedback [700]. For example, ORPS [790] and Drori et al. [129] minimize dependence on human annotations by using execution feedback, enabling models to autonomously refine their solutions. Additionally, Shrestha et al. [531] contribute by translating model outputs into Python code, helping to identify logical errors, gain insights into flawed reasoning processes, and guide improvements in mathematical reasoning. Xu et al. [711] integrate reasoning models with an interactive environment, enabling learning in more dynamic scenarios and creating a more generalizable self-learning framework.